HUBBLE SPACE TELESCOPE DAILY REPORT #1096 PERIOD COVERED: 0700 EST Daily Status Report as of 075/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 5090 (Large Area/Multicolor Survey: Cycle 4) The WF/PC-2 was used to make deep exposure images of fields #1415, #1416, and #1418 as part of a large area, high galactic latitude, multicolor sky survey to learn more about the distribution of galaxies, their morphology, distribution in magnitude and redshift, and evolution of galactic luminosity, function, and colors. A total 19 images were received and displayed. Some recentering events were reported to have occurred, but they did not appear to degrade the data. 1.2 Completed GHRS/2 5549 (GHRS/2 Side 2 Dark Noise Monitor: Cycle 4) The High Resolution Spectrograph (Side One) was used to make internal tests for routine monitoring of detector dark noise. There was no carousel motion. A total of sixteen observations were received and displayed. They all were reported to be nominal. 1.3 Completed GHRS/2 5546 (GHRS Side 2 Threshold Adjustment Test: Cycle 4) The High Resolution Spectrograph was used to conduct an engineering test to make internal threshold level measurements for the noisy channels. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were noted. 1.4 Completed Two Sets of 5570 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. A total of four images were received and displayed. One image was underexposed. Additional data is scheduled to be downlinked from the science tape recorder later. 1.5 Completed WF/PC-2 5476 (Continuum Snapshots of 3RC Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C277.3 as part of a continuum survey and to study the galaxy's morphology. Three images were received and displayed. The observations were reported to look okay. Additional data still remains on the science tape recorder. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled :8 Successful :8 2.3 Significant HSTARs: None 2.4 Operations Notes: Another nominal day was experienced as all SMOV activities reported no operational problems, and all HST subsystems performed flawlessly. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WF/PC-2 5090 (Large Area/Multicolor Survey: Cycle 4) 3.2 Complete WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Large Set) 3.3 Complete GHRS/2 5549 (GHRS Side 2 Dark Noise Monitor: Cycle 4) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-